# Convergence and long-time behavior of finite volumes for a generalized Poisson-Nernst-Planck system with cross-diffusion and size exclusion

### v1

@@ 2.1. Space and time discretizations
Let $\mathcal{T}$ denote a family of non-empty, disjoint, convex, open and polygonal control volumes $K\in\mathcal{T}$, whose Lebesgue measure is denoted by $m_{K}$. We also assume that control volumes partition the domain, meaning that $\overline{\Omega}=\bigcup_{K\in\mathcal{T}}\overline{K}$. Further, we call $\mathcal{E}$ a family of edges/faces, where $\sigma\in\mathcal{E}$ is a closed subset of $\overline{\Omega}$ contained in a hyperplane of $\mathbb{R}^{d}$. Each $\sigma$ has a strictly positive $(d-1)$-dimensional Hausdorff (or Lebesgue) measure, denoted by $m_{\sigma}$. We use the abbreviation $K|L=\partial K\cap\partial L$ for the intersection between two distinct control volumes which is either empty or reduces to a face contained in $\mathcal{E}$. The subset of all interior faces is denoted by
$\mathcal{E}_{\text{int}}=\{\sigma\in\mathcal{E}\text{ s.t. }\sigma=K|L\text{ for some }K,L\in\mathcal{T}\}.$
For any $K\in\mathcal{T}$, we assume that there exists a subset $\mathcal{E}_{K}$ of distinct elements of $\mathcal{E}$ such that the boundary of a control volume can be described by $\partial K=\bigcup_{\sigma\in\mathcal{E}_{K}}\sigma$ and, consequently, it follows that $\mathcal{E}=\bigcup_{K\in\mathcal{T}}\mathcal{E}_{K}$. Additionally, we assume that boundary edges $\mathcal{E}_{\text{ext}}=\mathcal{E}\setminus\mathcal{E}_{\text{int}}$ are either subsets of $\Gamma^{D}$ or $\Gamma^{N}$. To each control volume $K\in\mathcal{T}$ we assign a cell center ${x}_{K}\in K$ which satisfies the orthogonality condition: If $K,L$ share a face $\sigma=K|L$, then the vector $\overline{{x}_{K}{x}_{L}}$ is orthogonal to $\sigma=K|L$. The triplet $\left(\mathcal{T},\mathcal{E},\{{x}_{K}\}_{K\in\mathcal{T}}\right)$ is called an admissible mesh.
We introduce the notation $d_{\sigma}$ for the Euclidean distance between ${x}_{K}$ and ${x}_{L}$ if $\sigma=K|L$ or between ${x}_{K}$ and the affine hyperplane spanned by $\sigma$ if $\sigma\subset\partial{\Omega}$. We also denote by $d_{K\sigma}=\operatorname{dist}(x_{K},\sigma)$, so that $d_{\sigma}=d_{K\sigma}+d_{L\sigma}$ if $\sigma=K|L\in\mathcal{E}_{\text{int}}$ and $d_{\sigma}=d_{K\sigma}$ if $\sigma\in\mathcal{E}_{K}\cap\mathcal{E}_{\text{ext}}$. The transmittivity of the edge $\sigma\in\mathcal{E}$ is defined by ${a_{\sigma}=\frac{m_{\sigma}}{d_{\sigma}}}$. The size of the mesh is
@@
For the time discretization, we decompose the time interval $\mathbb{R}_{+}:=[0,+\infty)$ into an unbounded increasing sequence $\left(t^{n}\right)_{n\geq 0}$ with $t^{0}=0$ and possibly non-uniform time steps
@@
We finally introduce $\Delta t=\sup_{n\in\mathbb{N}\setminus\{0\}}\tau^{n}$, which we assume to be finite.
@@ 2.2. The finite volume schemes
$\phi_{K\sigma}^{n}=\begin{cases}\phi_{L}^{n}&\text{if}\;\sigma=K|L\in\mathcal{E}_{\text{int}},\\ {\phi_{K}^{n}}&{\text{if}\;\sigma\subset\Gamma^{N},}\\ \phi_{\sigma}^{D}=\frac{1}{m_{\sigma}}\int_{\sigma}\phi^{D}&\text{if}\;\sigma\subset\Gamma^{D}.\end{cases}$
@@
We would like to emphasise that the construction of the scheme which we will refer to as the SG scheme is not based on the original idea of [36]. Rather, we take advantage of the free-energy diminishing character of the SG scheme highlighted in [18] and exploited in [4, 37].
@@
(30) $\sum_{i=0}^{I}u_{i,K}^{0}=1\;\text{for all $K\in\mathcal{T}$},\;\text{and}\;\sum_{K\in\mathcal{T}}u_{i,K}^{0}m_{K}=\int_{\Omega}u_{i}^{0}>0\quad\text{for $i=0,\dots,I$}.$
@@
2.3. Main results and organisation of the paper
We provide here simple presentations of our main results, which will be detailed and proven in the next sections. As already mentioned, the goals of this contribution is twofold. It first aims to complement the elements of analysis sketched in the contribution [11], by showing existence of discrete solutions and convergence of the scheme. Secondly it intends to investigate the long time behavior of the scheme.
@@ 2.3. Main results and organisation of the paper
${0<u_{i,K}^{n}<1}\qquad\forall i=0,\dots,I,\;K\in\mathcal{T},\;n\geq 1$
@@
Theorem 2.1 states that, at fixed mesh and time step, the numerical scheme admits a positive solution and preserves the $L^{\infty}$-bounds. Moreover, the inequality (32) is a discrete counterpart to (9). The discrete version $\mathcal{H}_{\mathcal{T}}^{n}$ of the free energy functional (11) is rigorously defined in formula (41). Schemes that verify (32) in addition to more classical properties like mass conservation or positivity preservation are often referred to in the literature as thermodynamically consistent or structure-preserving. Our scheme, therefore, falls into this class of schemes.
Theorem 2.1 allows to define the so-called approximate solution to the problem, which consists of piecewise constants functions $U_{\mathcal{T},\tau}=\left(u_{i,\mathcal{T},\tau}\right)_{1\leq i\leq I}$ and $\phi_{\mathcal{T},\tau}$ defined by
@@
$\phi_{\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}\phi\quad\text{and}\quad u_{0,\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}u_{0}\quad\text{in}\;L^{p}_{\text{loc}}(\mathbb{R}_{+};L^{p}(\Omega))\quad\forall p\in[1,+\infty),$
@@
[…]
@@
Our last main theoretical result is about the long time asymptotic of the scheme for a fixed mesh and sequence of time steps. We denote by $\Psi_{\mathcal{T}}:\mathbb{R}^{\mathcal{T}}\times\mathbb{R}^{I}\to\mathbb{R}$ the functional defined by
@@
${y}_{K\sigma}=\begin{cases}{y}_{L}&\text{if}\;\sigma=K|L\in\mathcal{E}_{\text{int}},\\ {y}_{K}&\text{if}\;\sigma\subset\Gamma^{N},\\ \phi_{\sigma}^{D}&\text{if}\;\sigma\subset\Gamma^{D}.\end{cases}$
@@
[…]
We deduce from this theorem an efficient way to compute the long-time behavior of the scheme, which consists in seeking for $\left(({\phi_{K}^{\infty})}_{K\in\mathcal{T}},{(\mu_{i,\mathcal{T}}^{\infty})}_{1\leq i\leq I}\right)$ belonging to $\operatorname{argmin}\Psi_{\mathcal{T}}$ and to a posteriori reconstruct the volume fraction profiles thanks to the functions $v_{i}$ defined by (18). We also remark that the uniqueness of the long time limit is established in the discrete setting, while the question remains open in the continuous setting. Conjecturing the uniqueness of such an asymptotic limit for the continuous model, which then has to be the unique minimizer of $\Psi$ defined by (23), then Theorem 2.3 illustrates in a yet another aspect the structure preserving nature of our scheme.
The rest of the paper is organized as follows. Section 3 is devoted to the proof of Theorem 2.1, as well as to some post-treatment of the energy / energy dissipation inequality (32) in order to derive some uniform bounds on quantities to be used in the convergence analysis carried out in Section 4, where the proof of Theorem 2.2 is detailed. Our last main theoretical result Theorem 2.3 is proved in Section 5. The numerical results presented in Section 6 contains some numerical evidence of the convergence of the scheme. They also illustrate that the situation is more complex than what Theorem 2.3 suggests at first glance. Although the approximate volume fractions are (strictly) positive as pointed out in Theorem 2.1, $u_{0,\mathcal{T},\tau}$ can take very small values in some parts of the domain. The discrete fluxes (26) then take very small values as well, making the evolution extremely slow. As the linear and nonlinear systems corresponding to the scheme are merely solved approximately, one furthermore has to pay a particular attention to preserve the positivity at the discrete level.
@@ 3. Uniform a priori bounds and existence of a discrete solution
The goal of this section is to prove Theorem 2.1, i.e. to show that the nonlinear system corresponding to the scheme (24)–(27) admits at least one solution, and that beyond local conservativity, this solution preserves at the discrete level some key features of the model, namely the positivity of the volume fractions, as well as uniform $L^{\infty}$ bounds for them, and the decay of the free energy. Furthermore, in this part, precise quantification of the dissipated total free energy supplies enough uniform estimates, in the mesh size $h_{\mathcal{T}}$ and time steps $\left(\tau_{n}\right)_{n\geq 1}$, to perform the convergence analysis in Section 4. The grid $\mathcal{T}$ and the time steps $\left(\tau^{n}\right)_{n\geq 1}$ remain fixed throughout this section.
@@ 3.1. Uniform a priori bounds and existence of a discrete solution
Since our scheme is locally conservative, i.e., $F_{K\sigma}^{n}+F_{L\sigma}^{n}=0$ for all $\sigma=K|L\in\mathcal{E}_{\text{int}}$, then summing (25) over $K$ shows by induction and thanks to (29) the result. ∎
@@
$0\geq u_{0,K}^{n}=u_{0,K}^{n-1}+\frac{\tau^{n}}{m_{K}}\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{K}}F_{i,K\sigma}^{n}\geq 0$
In particular, all the fluxes $F_{i,K\sigma}^{n}$, $i=1,\dots,I$ and $\sigma\in\mathcal{E}_{K}$ are equal to 00. In view of formula (37) and of the strict positivity of $\mathfrak{B}$, this implies either that $u_{i,K}^{n}\leq 0$ for all $i$, which yields a contradiction with (27), or that $u_{0,L}^{n}\leq 0$ for all the cells $L$ sharing an edge $\sigma=K|L$ with $K$. Since $\Omega$ is connected, one would obtain that $u_{0,K}^{n}=0$ for all $K\in\mathcal{T}$ and thus that $\sum_{K\in\mathcal{T}}u_{0,K}^{n}m_{K}=0$. This contradicts (36), and thus we necessarily have that $u_{0,K}^{n}>0$ for all $K\in\mathcal{T}$.
@@
$0\geq u_{i,K}^{n}=u_{i,K}^{n-1}-\frac{\tau^{n}}{m_{K}}\sum_{\sigma\in\mathcal{E}_{K}}F_{i,K\sigma}^{n}\geq 0$
This leads to $u_{i,K}^{n}=0$ and to $F_{i,K\sigma}^{n}=0$ for all $\sigma\in\mathcal{E}_{K}$. Since we already know that $u_{0,K}^{n}>0$, we deduce from (37) that $u_{i,L}^{n}\leq 0$ for all cell $L$ sharing a cell $\sigma=K|L$ with $K$. As above, this implies as $u_{0,K}^{n}=0$ for all $K\in\mathcal{T}$, which contradicts (36). Then $u_{i,K}^{n}>0$ for all $K\in\mathcal{T}$, concluding the proof of Proposition 3.2. ∎
A consequence of previous proposition is that a solution to the modified scheme with (37) instead of (26) is also a solution to the original scheme (24)–(27).
@@
Proposition 3.3.
@@
The proof is based on an inductive procedure in time. At each time step $n\geq 1$, we use a topological degree argument to show the existence of a solution $(U^{n}_{K},\phi^{n}_{K})_{K\in\mathcal{T}}\in(0,1)^{I\times\mathcal{T}}\times[-C,C]^{\mathcal{T}}$ to (24)-(25), with the fluxes defined as in (37). In fact, the Proposition 3.2, ensures the existence for the scheme (24)-(26).
@@
:$\mathcal{H}\;:\;\begin{cases}[0,\tau^{n}]\times[-2,2]^{I\times\mathcal{T}}\times[-C-1,C+1]^{\mathcal{T}}\to\mathbb{R}^{\mathcal{T}}\times\mathbb{R}^{\mathcal{T}}\\ (s,U,\phi)\mapsto\mathcal{H}(s,U,\phi)\end{cases}$
whose zeros $(s,U^{(s)},\phi^{(s)})$ are the solutions of (40), is hence an homotopy. Furthermore, all along it, its zeros $(U^{(s)},\phi^{(s)})$ remain inside the compact subset $[0,1]^{I\times\mathcal{T}}\times[-C,C]^{\mathcal{T}}$.
Thus, the topological degree corresponding to $\mathcal{H}(s,U,\phi)=0$ is equal to one, all along the homotopy, and hence in particular for $s=\tau^{n}$. That implies the existence of a solution to the scheme (but it does not guarantee uniqueness). ∎
We note that the proof proposed here is simpler than that found in, for example,  [14, Proposition 3.2]. The reason lies in the fact that we did not use entropic inequality to derive the uniform estimates on volume fractions and, therefore, we do not have to guarantee the validity of it during the homotopy.

### v2

@@ 2.1. Space and time discretizations
Let $\mathcal{T}$ denote a family of non-empty, disjoint, convex, open, and polygonal control volumes $K\in\mathcal{T}$, whose Lebesgue measure is denoted by $m_{K}$. We also assume that control volumes partition the domain, meaning that $\overline{\Omega}=\bigcup_{K\in\mathcal{T}}\overline{K}$. Further, we call $\mathcal{E}$ a family of edges/faces, where each $\sigma\in\mathcal{E}$ is a closed subset of $\overline{\Omega}$ contained in a hyperplane of $\mathbb{R}^{d}$. Each $\sigma$ has a strictly positive (d1)(d-1)-dimensional Hausdorff (or Lebesgue) measure, denoted by $m_{\sigma}$. We use the abbreviation $K|L=\partial K\cap\partial L$ for the intersection between two distinct control volumes which is either empty or reduces to a face contained in $\mathcal{E}$. The subset of all interior faces is denoted by
$\mathcal{E}_{\rm int}=\left\{\sigma\in\mathcal{E}\;\middle|\;\sigma=K|L\text{ for some }K,L\in\mathcal{T}\right\}.$
For any $K\in\mathcal{T}$, we assume that there exists a subset $\mathcal{E}_{K}$ of distinct elements of $\mathcal{E}$ such that the boundary of a control volume can be described by $\partial K=\bigcup_{\sigma\in\mathcal{E}_{K}}\sigma$ and, consequently, $\mathcal{E}=\bigcup_{K\in\mathcal{T}}\mathcal{E}_{K}$. Additionally, we assume that boundary edges $\mathcal{E}_{\text{ext}}=\mathcal{E}\setminus\mathcal{E}_{\rm int}$ are either subsets of $\Gamma^{D}$ or $\Gamma^{N}$. To each control volume $K\in\mathcal{T}$ we assign a cell center ${x}_{K}\in K$ which satisfies the orthogonality condition: If $K,L$ share a face $\sigma=K|L$, then the vector $x_{L}-x_{K}$ is orthogonal to $\sigma=K|L$.The triplet $\left(\mathcal{T},\mathcal{E},\{{x}_{K}\}_{K\in\mathcal{T}}\right)$ is called an admissible mesh.
A classical way to construct such admissible mesh is to place some seeds $(x_{K})_{K\in\mathcal{T}}$ in the domain $\Omega$ and to define the corresponding Voronoï  tessellation. As in practice the condition $x_{K}\in K$ can be relaxed, but sill with the constraint $x_{L}-x_{K}$ points outward with respect to $K$, another popular choice to build admissible meshes from a simplicial mesh satisfying the Delaunay condition is to choose $x_{K}$ as the center of the circumcircle of the simplex $K$.
We introduce the notation $d_{\sigma}$ for the Euclidean distance between ${x}_{K}$ and ${x}_{L}$ if $\sigma=K|L$ or between ${x}_{K}$ and the affine hyperplane spanned by $\sigma$ if $\sigma\subset\partial{\Omega}$. We also denote by $d_{K\sigma}=\operatorname{dist}(x_{K},\sigma)$, so that $d_{\sigma}=d_{K\sigma}+d_{L\sigma}$ if $\sigma=K|L\in\mathcal{E}_{\rm int}$ and $d_{\sigma}=d_{K\sigma}$ if $\sigma\in\mathcal{E}_{K}\cap\mathcal{E}_{\text{ext}}$. The transmittivity of the edge $\sigma\in\mathcal{E}$ is defined by $a_{\sigma}=\frac{m_{\sigma}}{d_{\sigma}}$}. The size of the mesh is
@@
For the time discretization, we decompose the time interval $\mathbb{R}_{\geq 0}:=[0,+\infty)$ into an unbounded increasing sequence $\left(t^{n}\right)_{n\geq 0}$ with $t^{0}=0$ and possibly non-uniform time steps
@@
[…]
@@ 2.2. The finite volume schemes
$\phi_{K\sigma}^{n}=\begin{cases}\phi_{L}^{n}&\text{if}\;\sigma=K|L\in\mathcal{E}_{\rm int},\\ {\phi_{K}^{n}}&{\text{if}\;\sigma\subset\Gamma^{N},}\\ \phi_{\sigma}^{D}=\frac{1}{m_{\sigma}}\int_{\sigma}\phi^{D}&\text{if}\;\sigma\subset\Gamma^{D}.\end{cases}$
@@
We would like to emphasize that the construction of the scheme which we will refer to as the SG scheme is not based on the original idea of [39]. Rather, we take advantage of the free-energy diminishing character of the SG scheme highlighted in [18] and exploited in [4, 40].
@@
[…]
@@
2.3. Main results and organization of the paper
We provide here simple presentations of our main results, which will be detailed and proven in the next sections. As already mentioned, the goals of this contribution are twofold. It first aims to complement the elements of analysis sketched in the contribution [11], by showing existence of discrete solutions and convergence of the scheme. Secondly it intends to investigate the long-time behavior of the scheme.
@@ 2.3. Main results and organization of the paper
$0<u_{i,K}^{n}<1\qquad\forall i=0,\dots,I,\;K\in\mathcal{T},\;n\geq 1.$
@@
Theorem 2.1 states that, at fixed mesh and time step, the numerical scheme admits a positive solution and preserves the $L^{\infty}$-bounds. Moreover, the inequality (32) is a discrete counterpart to (9). The discrete version $\mathcal{H}_{\mathcal{T}}^{n}$ of the free energy functional (11) is rigorously defined in formula (41). Schemes that verify (32) in addition to more classical properties such as mass conservation or positivity preservation are often referred to in the literature as thermodynamically consistent or structure-preserving. Our scheme, therefore, falls into this class of schemes.
Theorem 2.1 allows to define the so-called approximate solution to the problem, which consists of functions $U_{\mathcal{T},\tau}=\left(u_{i,\mathcal{T},\tau}\right)_{1\leq i\leq I}$ and $\phi_{\mathcal{T},\tau}$, piecewise constant in both space and time, defined by
@@
$\phi_{\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}\phi\quad\text{and}\quad u_{0,\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}u_{0}\quad\text{in}\;L^{p}_{\rm loc}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}};L^{p}(\Omega))\quad\forall p\in[1,+\infty),$
@@
$U_{\mathcal{T}_{\ell},\tau_{\ell}}\underset{\ell\to+\infty}{\longrightarrow}U\quad\text{in the }L^{\infty}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathbb{R}_{\geq 0}}\times\Omega)^{I}\text{ weak-$\star$ sense}.$
@@
Our last main theoretical result is about the long-time asymptotic of the scheme for a fixed mesh and sequence of time steps. We denote by $\Psi_{\mathcal{T}}:\mathbb{R}^{\mathcal{T}}\times\mathbb{R}^{I}\to\mathbb{R}$ the functional defined by
@@
${y}_{K\sigma}=\begin{cases}{y}_{L}&\text{if}\;\sigma=K|L\in\mathcal{E}_{\rm int},\\ {y}_{K}&\text{if}\;\sigma\subset\Gamma^{N},\\ \phi_{\sigma}^{D}&\text{if}\;\sigma\subset\Gamma^{D}.\end{cases}$
@@
$u_{i,K}^{n}\underset{n\to+\infty}{\longrightarrow}v_{i}(\phi_{K}^{\infty},\boldsymbol{\mu}_{\mathcal{T}}^{\infty})\qquad\forall K\in\mathcal{T},\;0\leq i\leq I.$
We deduce from this theorem an efficient way to compute the long-time behavior of the scheme, which consists in seeking for $\left(({\phi_{K}^{\infty})}_{K\in\mathcal{T}},{(\mu_{i,\mathcal{T}}^{\infty})}_{1\leq i\leq I}\right)$ belonging to $\operatorname{argmin}\Psi_{\mathcal{T}}$ and to a posteriori reconstruct the volume fraction profiles thanks to the functions $v_{i}$ defined by (18). We also remark that the uniqueness of the long-time limit is established in the discrete setting, while the question remains open in the continuous setting. Conjecturing the uniqueness of such an asymptotic limit for the continuous model, which then has to be the unique minimizer of $\Psi$ defined by (23), then Theorem 2.3 illustrates in yet another aspect the structure-preserving nature of our scheme.
The rest of the paper is organized as follows. Section 3 is devoted to the proof of Theorem 2.1, as well as to some post-treatment of the energy / energy dissipation inequality (32) in order to derive some uniform bounds on quantities to be used in the convergence analysis carried out in Section 4, where the proof of Theorem 2.2 is detailed. Our last main theoretical result Theorem 2.3 is proved in Section 5. The numerical results presented in Section 6 contain some numerical evidence of the convergence of the scheme. They also illustrate that the situation is more complex than what Theorem 2.3 suggests at first glance. Although the approximate volume fractions are (strictly) positive as pointed out in Theorem 2.1, $u_{0,\mathcal{T},\tau}$ can take very small values in some parts of the domain. The discrete fluxes (26) then take very small values as well, making the evolution extremely slow. As the linear and non-linear systems corresponding to the scheme are merely solved approximately, one furthermore has to pay a particular attention to preserve the positivity at the discrete level.
@@ 3. Uniform a priori bounds and existence of a discrete solution
The goal of this section is to prove Theorem 2.1, i.e. to show that the non-linear system corresponding to the scheme (24)–(27) admits at least one solution, and that beyond local conservativity, this solution preserves at the discrete level some key features of the model, namely the positivity of the volume fractions, as well as uniform $L^{\infty}$ bounds for them, and the decay of the free energy. Furthermore, in this part, precise quantification of the dissipated total free energy supplies enough uniform estimates, in the mesh size $h_{\mathcal{T}}$ and time steps $\left(\tau_{n}\right)_{n\geq 1}$, to perform the convergence analysis in Section 4. The grid $\mathcal{T}$ and the time steps $\left(\tau^{n}\right)_{n\geq 1}$ remain fixed throughout this section.
@@ 3.1. Uniform a priori bounds and existence of a discrete solution
Since our scheme is locally conservative, i.e., $F_{i,K\sigma}^{n}+F_{i,L\sigma}^{n}=0$ for all $\sigma=K|L\in\mathcal{E}_{\rm int}$, $i=1,\dots,I$, then summing (25) over $K$ shows by induction and thanks to (29) the result. ∎
@@
$0\geq u_{0,K}^{n}=u_{0,K}^{n-1}+\frac{\tau^{n}}{m_{K}}\sum_{i=1}^{I}\sum_{\sigma\in\mathcal{E}_{K}}F_{i,K\sigma}^{n}\geq 0.$
In particular, all the fluxes $F_{i,K\sigma}^{n}$, $i=1,\dots,I$ and $\sigma\in\mathcal{E}_{K}$, are equal to 0. In view of formula (37) and of the strict positivity of $\mathfrak{B}$, this implies either that $u_{i,K}^{n}\leq 0$ for all $i$, which yields a contradiction with (27), or that $u_{0,L}^{n}\leq 0$ for all the cells $L$ sharing an edge $\sigma=K|L$ with $K$. Since $\Omega$ is connected, one would obtain that $u_{0,K}^{n}=0$ for all $K\in\mathcal{T}$ and thus that $\sum_{K\in\mathcal{T}}u_{0,K}^{n}m_{K}=0$. This contradicts (36), and thus we necessarily have that $u_{0,K}^{n}>0$ for all $K\in\mathcal{T}$.
@@
$0\geq u_{i,K}^{n}=u_{i,K}^{n-1}-\frac{\tau^{n}}{m_{K}}\sum_{\sigma\in\mathcal{E}_{K}}F_{i,K\sigma}^{n}\geq 0.$
This leads to $u_{i,K}^{n}=0$ and to $F_{i,K\sigma}^{n}=0$ for all $\sigma\in\mathcal{E}_{K}$. Since we already know that $u_{0,K}^{n}>0$, we deduce from (37) that $u_{i,L}^{n}\leq 0$ for all cells $L$ sharing a face $\sigma=K|L$ with $K$. As above, this implies as $u_{0,K}^{n}=0$ for all $K\in\mathcal{T}$, which contradicts (36). Then $u_{i,K}^{n}>0$ for all $K\in\mathcal{T}$, concluding the proof of Proposition 3.2. ∎
A consequence of the previous proposition is that a solution to the modified scheme with (37) instead of (26) is also a solution to the original scheme (24)–(27).
@@
Proposition 3.3 (Uniform bound for the electric potential).
@@
The proof is based on an inductive procedure in time. At each time step $n\geq 1$, we use a topological degree argument to show the existence of a solution $(U^{n}_{K},\phi^{n}_{K})_{K\in\mathcal{T}}\in(0,1)^{I\times\mathcal{T}}\times[-C,C]^{\mathcal{T}}$ to (24)–(25), with the fluxes defined as in (37). Because of the positivity of the discrete volume fractions stated in Proposition 3.2 and to the equivalence between (37) and (26) for such positive volume fractions, this ensures the existence for the original scheme (24)–(26).
@@
${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}}\;:\;\begin{cases}[0,\tau^{n}]\times[-2,2]^{I\times\mathcal{T}}\times[-C-1,C+1]^{\mathcal{T}}\to\mathbb{R}^{\mathcal{T}}\times\mathbb{R}^{\mathcal{T}}\\ (s,U,\phi)\mapsto{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}}(s,U,\phi)\end{cases}$
whose zeros $(s,U^{(s)},\phi^{(s)})$ are the solutions of (40), is hence an homotopy. Furthermore, all along it, its zeros $(U^{(s)},\phi^{(s)})$ remain inside the compact subset $0,1]^{I\times\mathcal{T}}\times[-C,C]^{\mathcal{T}}$.
Thus, the topological degree corresponding to ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathcal{S}}(s,U,\phi)=0$ is equal to one, all along the homotopy, and hence in particular for $s=\tau^{n}$. That implies the existence of a solution to the scheme (but it does not guarantee uniqueness). ∎
We note that the proof proposed here is simpler than that found in, for example,  [14, Proposition 3.2]. The reason lies in the fact that we did not use an entropy inequality to derive the uniform estimates on volume fractions and, therefore, we do not have to guarantee the validity of it during the homotopy.
Remark 3.5 (On the uniqueness of solutions).
The uniqueness of weak solutions for (1)–(7), as well as for numerical schemes found in the literature, remains largely unresolved. This challenge is inherent to the cross-diffusion structure of the problem. In [31], uniqueness is established under strong assumptions on the diffusion coefficients and valences, specifically $D_{i}\equiv D$ and $z_{i}\equiv z$ for all $i=1,\dots,I$.
For the numerical scheme (25)–(29), by following the argument in [7, Section 3.2, Uniqueness of solutions], uniqueness of discrete solutions can also be shown under the same assumptions: constant diffusion coefficients $D_{i}\equiv D$ and the absence of a drift term in the fluxes. The assumption of constant diffusion coefficients enables a preliminary proof of uniqueness for the discrete variable $u_{0}$. In our scheme, since no drift term is present and $\mathfrak{B}(0)=1$, the function $\mathfrak{B}$ does not appear in the definition of the fluxes. Consequently, the equation for the discrete $u_{0}$ is significantly simpler than that in [7], and uniqueness of $u_{0}$ follows straightforwardly. The absence of a drift term also allows the application of Gajewski’s entropy method to deduce uniqueness for the remaining components $(u_{i})_{1\leq i\leq I}$, as in [7].